OpenStudy (anonymous):

Simplify the expression 8/2*4

OpenStudy (uglyturtle):

8/2 = 4 4 * 4 = 16

rvc (rvc):

(8/2)^4 is it so?

OpenStudy (anonymous):

so it's 4/16 ?

OpenStudy (nirmalnema):

8/8 = 1

rvc (rvc):

8*8*8*8/2*2*2*2=?

OpenStudy (anonymous):

the answer is 1 ?

rvc (rvc):

8/2*2*2*2 = 8/8 =1

rvc (rvc):

@cancherolugano can u tell me the ans of given question?

rvc (rvc):

ans me